OpenStudy (andrewkaiser333):

You increase your walking speed from 1 m/s to 3 m/s in a period of 1 s. What is your acceleration? 4 m/s2 3 m/s2 5 m/s2 2 m/s2 I think the answer is A

OpenStudy (andrewkaiser333):

@MysticFury

OpenStudy (michele_laino):

the magnitude of the requested acceleration, is: \[\Large a = \frac{{{v_2} - {v_1}}}{{\Delta t}} = \frac{{3 - 1}}{1} = ...?\]

OpenStudy (andrewkaiser333):

so D?

OpenStudy (michele_laino):

yes!

OpenStudy (andrewkaiser333):

thanks!!!